### v1

# M2CVD: Multi-Model Collaboration for Code Vulnerability Detection

Abstract.
Large Language Models (LLMs) have strong capabilities in code comprehension, but fine-tuning costs and semantic alignment issues limit their project-specific optimization; conversely, code models such CodeBERT are easy to fine-tune, but it is often difficult to learn vulnerability semantics from complex code languages. To address these challenges, this paper introduces the Multi-Model Collaborative Vulnerability Detection approach (M2CVD) that leverages the strong capability of analyzing vulnerability semantics from LLMs to improve the detection accuracy of code models. M2CVD employs a novel collaborative process: first enhancing the quality of vulnerability semantic description produced by LLMs through the understanding of project code by code models, and then using these improved vulnerability semantic description to boost the detection accuracy of code models. We demonstrated M2CVD’s effectiveness on two real-world datasets, where M2CVD significantly outperformed the baseline. In addition, we demonstrate that the M2CVD collaborative method can extend to other different LLMs and code models to improve their accuracy in vulnerability detection tasks.
Vulnerability detection, Model collaboration, Large language model, Pre-trained models
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Vulnerabilities in software refer to code weaknesses that can be easily exploited, which can lead to serious consequences such as unauthorized information disclosure  (Fu and Tantithamthavorn, 2022) and cyber extortion  (Thapa et al., 2022). Recent statistics underscore this burgeoning issue: In Q1 of 2022, the US National Vulnerability Database (NVD) disclosed 8,051 vulnerabilities, marking a 25% surge from the previous year  (Cheng et al., 2022). Further accentuating this trend, a study revealed that out of 2,409 analyzed codebases, 81% had at least one recognized open-source vulnerability. The increasing scale and ubiquity of these vulnerabilities emphasize the need for well-developed automated vulnerability detection mechanisms. Such a detection system helps to strengthen software security and forestall a range of potential security risks (Fu and Tantithamthavorn, 2022; Thapa et al., 2022; Jang-Jaccard and Nepal, 2014; Johnson et al., 2011).
The vulnerability detection models in the existing literature are mainly divided into two categories: (1) conventional detection models  (Yamaguchi, 2015, 2017) and (2) Deep Learning(DL)-based models  (Li et al., 2021a; Duan et al., 2019; Russell et al., 2018; Li et al., 2018; Dam et al., 2017) The former typically requires experts to manually formulate detection rules (Checkmarx, 2022; Flawfinder, 2022). These methods are usually labor-intensive to create and are difficult to achieve low false positive rates and low false negative rates  (Li et al., 2018, 2021b). On the contrary, deep learning (DL) -based detection methods learn the patterns of vulnerabilities from a training set (Li et al., 2018; Dam et al., 2017; Lin et al., 2017). They avoid manual heuristic methods and autonomously learn and identify vulnerability features. In order to further learn the semantics of vulnerabilities, methods based on pre-trained code models (Nguyen et al., 2022) and vulnerability detection studies based on LLMs (Fu et al., 2023) have been proposed successively. In summary, traditional vulnerability detection methods are usually depend on pre-defined rules, a process of expert interventions, rendering them laborious and occasionally imprecise. The DL-based detection methods, by contrast, show better detection ability by learning vulnerable code patterns automatically.
In the latest research (Steenhoek et al., 2023a), the efficacy of pre-trained language models for software vulnerability detection has been extensively explored, encompassing LLMs such as ChatGPT (openAI, 2022) and LLaMa (Llama, 2022), alongside pre-trained code models like CodeBERT (Feng et al., 2020) and UniXcoder (Guo et al., 2022). Compared with traditional deep learning networks, these models show more excellent performance in code vulnerability detection tasks after fine-tuning (Nguyen et al., 2022).
Detecting vulnerabilities using pre-trained models has its benefits, but when it comes to real-world applications, we encounter the first challenge that the complexity of code makes it hard for the code model to learn vulnerability semantics (Steenhoek et al., 2023b). The pre-training datasets for code models usually do not have vulnerability descriptions. Though code models can be fine-tuned on a domain-specific vulnerability dataset, these datasets usually only contain labels to show if a piece of code is vulnerable or not. Without semantic description, it would be difficult for code models to learn the actual cause of the vulnerabilities. For the same reason, existing vulnerability detection methods usually only output a vulnerability judgment indicating whether the code is vulnerable.
In contrast to existing approaches that use pattern matching to enhance vulnerability semantic  (Steenhoek et al., 2023b), we resort tothe strong understanding abilities of LLMs to create natural language descriptions of vulnerable code, so as to make connections between code and the causes of vulnerabilities. This will bring two benefits: the more abstract natural language description will help the code model better learn the semantics of the vulnerability, and the vulnerability description can help programmers better determine the cause of the vulnerability to maintain the code. But we’ve hit the second challenge, the semantic alignment problem of LLMs. Given the scale of LLMs, fine-tuning them on a specific domain is tough. Real projects, organizations, or specific fields have their own coding rules and business logic. Using LLMs trained on data from open domain might not make accurate vulnerability judgment on code in a specific domain (e.g., Fu et al. (2023) reports a F1 score of 29% with GPT4). As a result, LLMs may generate incorrect vulnerability description.
In this paper, we’ve proposed M2CVD, an innovative approach that combines the strengths of pre-trained code models and LLMs to better detect vulnerabilities. For the first challenge, we rely on the ability of LLMs to interpret vulnerabilities, leveraging the explanatory text to hellp code models to understand the semantics of vulnerabilities. For the second challenge, we rely on the advantage of code models that is easy to fine-tune, and use their judgment results to enhance the vulnerability semantic understanding of LLMs for specific projects. In this way, M2CVD can help operators to improve the accuracy of vulnerability detection through the collaborative interaction process combined with LLMs API without changing the existing code model structure. In summary, the main contributions of this paper are as follows:
  • a)
    We propose M2CVD, an approach that integrates the capabilities of pre-trained code models and LLMs to better utilize their strengths for enhancing the precision of vulnerability detection tasks. Compared with the existing vulnerability detection, M2CVD supports the output of vulnerability semantic description to assist programmers to maintain code.
  • b)
    This paper proposes a vulnerability semantic description refinement method, which leverages the insights of fine-tuning pre-trained code models on specific data to effectively enhance the vulnerability description generation ability of unfine-tuned LLMs on project-specific domain code.
  • c)
    We evaluate our approach through extensive experimentation on two real-world datasets. Experimental results show that the M2CVD can still improve the performance of code vulnerability detection with the different of pre-trained code model and LLMs.  [^1] Our anonymous replication package (data and code) :https://figshare.com/articles/dataset/M2CVD/24831567.
2. Related work
2.1. Traditional Vulnerability Detection
Over the years, a lot of methods for vulnerability detection have emerged. Overall, initial research in this area focused on identifying vulnerabilities by means of manually customized rules  (Checkmarx, 2022; Flawfinder, 2022). While these approaches provide heuristic approaches to vulnerability detection, they require extensive manual analysis and formulation of defect patterns. In addition, syntactic elements are repeated in different code fragments, as prescribed by certain rules, have been observed to induce elevated rates of both false positives and false negatives  (Yamaguchi, 2015, 2017; Li et al., 2021b).
2.2. Deep Neural Network for Vulnerability Detection
To minimize human intervention, recent works have turned to employing neural network-based models for the extraction of vulnerability features from code fragments  (Dam et al., 2017; Russell et al., 2018). Existing deep learning-based vulnerability detection models predominantly bifurcate into two classifications: token-based and graph-based models.
Token-based models treat code as a linear sequence and use neural networks to learn vulnerability features from known cases, aiming to identify previously undetected vulnerabilities.  (Russell et al., 2018; Li et al., 2018; Cheng et al., 2021). For instance, Russell et al. harnessed the power of both recurrent neural networks (RNNs) and convolutional neural networks (CNNs) to learn feature sets from code token sequences tailored for vulnerability identification (Russell et al., 2018). Concurrently, Li et al.  (Li et al., 2018) employed BiLSTM  (Schuster and Paliwal, 1997) to encode a segmented version of input code, known as ’code gadget’, centered on key markers, especially library/API function calls. However, these token-based models often ignore the complexity of the source code structure, which may lead to inaccurate detection.
While focusing on token-based models, another research direction is to reveal the potential of graph-based methods in the field of vulnerability detection (Li et al., 2021a; Chakraborty et al., 2021; Zheng et al., 2021; Nguyen et al., 2022). DeepWukong  (Cheng et al., 2021) utilizes GNN for feature learning, which focuses on compressing code fragments into a dense, low-dimensional vector space to enhance the detection of a large number of vulnerability types. DeepTective  (Rabheru et al., 2021) confronts vulnerabilities common to PHP scripts such as SQL injection, Cross-Site Scripting (XSS), and command injection by deploying a combination of Gated Recurrent Units (GRUs) and Graph Convolutional Networks.
Graph-based detection models learn code structure through varied graph representations, utilizing neural networks for vulnerability detection (Wu et al., 2022; Cao et al., 2022). For instance, Zhou et al.  (Zhou et al., 2019) used the gated graph recurrent network  (Li et al., 2016), extracting structural details from triadic graph representations—AST, CFG, and DFG. Chakraborty et al.  (Chakraborty et al., 2021) introduced REVEAL, an innovative approach that amalgamation of the gated graph neural network, re-sampling techniques  (Chawla et al., 2002), and triplet loss  (Mao et al., 2019). Wu et al.  (Wu et al., 2022) proposed a approach that can efficiently convert the source code of a function into an image while preserving the program details. Meanwhile, Cao et al.  (Cao et al., 2022) proposed a statement-centric approach, based on flow-sensitive graph neural network, to understanding semantic and structural data.
2.3. Pre-Trained Models for Vulnerability Detection
Taking inspiration from the success of pre-trained models in the field of natural language processing (NLP), an increase of related research works aims to leverage these pre-trained models to improve code vulnerability detection accuracy  (Feng et al., 2020; Kanade et al., 2020; Niu et al., 2022; Lin et al., 2021; Bai et al., 2021; Guo et al., 2022).
The core idea of these works is a pre-trained model on a large amount of source code data, followed by specialized fine-tuning for a specific task.  (Kanade et al., 2020). To illustrate, Feng et al.  (Feng et al., 2020) proposed CodeBERT specifically for understanding and generating source code, which combines the processing power of natural and programming languages. Similarly, CuBERT combines masked language modeling with sentence prediction for code representation  (Kanade et al., 2020). In addition, some pre-trained models also take the structural information of the code fragment into account in the initial training phase (Niu et al., 2022; Lin et al., 2021). For example, Guo et al.’s GraphCodeBERT  (Hin et al., 2022) to infer different in the data flow of code fragments with the help of graph structures. DOBF  (Lachaux et al., 2021) introduces a novel pre-training objective predicated, explore whether such pre-training can enhance the model’s ability to learn the syntactic and structural complexity of source code. The objective is specifically tailored to address the structural dimension of programming languages. Concurrently, to enhance the graph-based representation, GraphCodeBERT  (Guo et al., 2020) proposed a pre-trained schema to seamlessly insert graph structure into Transformer-based architectures. This is achieved through the innovative use of graph-guided masked attention mechanisms, designed to mitigate noise in the data. The comparative evaluation positions CodeBERT as a baseline standard, which is a very classic pre-trained code model in a series of code-related tasks, including code clone detection and code translation. At the same time, UniXcoder as the latest pre-trained code model will also be done as a baseline method. UniXcoder, a unified cross-modal pre-trained programming language model, is trained on a large amount of code data as well as natural language data (Guo et al., 2022).
Since these pre-trained models have shown superior performance in various code-related tasks, some studies have attempted to use these models for vulnerability detection  (Fu and Tantithamthavorn, 2022; Thapa et al., 2022; Hin et al., 2022). However, if these models are directly used for vulnerability detection after fine-tuning with code data, they face the challenge of capturing vulnerability features from long code and complex structure (Zhang et al., 2023). Moreover, due to the nature of code data, the lack of vulnerability semantics information prevents these multi-modal models from taking full advantage of them. Therefore, we try to supplement the vulnerability semantics in the existing code data to reduce the cost of modeling and searching vulnerability features in complex code data.
Figure 1. The framework of M2CVD, which mainly contains three phase: 1. Initial vulnerability detection; 2. vulnerability description refinement and 3. Integrated vulnerability detection. The detection model uses historical vulnerability data to fine-tune, and then fine-tunes the validation model after the historical vulnerability data is supplemented by vulnerability semantics in phase 1 and phase 2.
3. Approach
In general, M2CVD requires three models to work together, including code model $f_{d}$,$f_{v}$ and LLM $f_{c}$, and relies on the collaborative interaction of $f_{d}$ and $f_{c}$ to assist the enhancement of $f_{v}$ model in the vulnerability detection task. The overall framework of M2CVD as shown in 1, consisting of three phases:
1) Phase I generates preliminary judgments and vulnerability descriptions with the help of $f_{c}$ and $f_{d}$.
2) In phase II, the judgments that are inconsistent with $f_{c}$ and $f_{d}$ in Phase I will be judged and described by $f_{c}$ for the second time.
3) The last phase uses the vulnerability text to enhance the vulnerability detection ability of $f_{v}$.
3.1. Initial Vulnerability Detection
We use $L=(P,Y)$ to represent the historical vulnerability dataset, where $p_{i}$ represents a code snippet in a programming language, and $y_{i}$ represents vulnerability labels, $0<i\leq M$. The $M$ is the number of code snippet. The values of $y_{i}$ are 00 or $1$. When $y_{i}=0$, it indicates that the code is free of vulnerabilities; conversely, it indicates a vulnerability in the code.
First, we split the vulnerability dataset $L$ into a training set $p_{t}$ and a validation set $p_{v}$. In the inference phase, the code to be detected is denoted as $p_{e}$. Then, we fine-tune the detection model using $p_{t}$. According to the methods provided by the existing literature (Steenhoek et al., 2023a), the detection model $f_{d}$ is obtained by fine-tuning on the historical vulnerability data.
After obtaining the detection model, the vulnerability assessment of the detection model and the vulnerability assessment and description of the LLM are obtained through the following two steps:
1) Generation of the assessment with detection model.
We need to use the detection model to complete the preliminary vulnerability assessment for $L$. The specific steps for this are:
(1) $z_{i}=f_{d}(p_{i}),0<i\leq M$
where $f_{d}$ represents the prediction step of the detection model, and $z_{i}$ denotes the detection model’s assessment of code snippet $p_{i}$. If the model determines that $p_{i}$ contains a vulnerability, then $z_{i}=0$; otherwise, $z_{i}=1$.
2) Generation of the assessment and description with LLMs. In this step, we conduct an initial vulnerability assessment and description of $p_{t},p_{v},p_{e}$ through an interactive approach using ChatGPT. We use the following prompt to obtain LLM’s assessment and description:
User:You are a senior programmer. Please evaluate the code below for vulnerabilities. If you believe there are vulnerabilities, reply starting with ’Yes’ and briefly explain the issue; otherwise, begin with ’No’. Code: int ff_get_wav_header(AVFormatContext *s, AVIOContext *pb ……
LLM: [Yes][Vulnerability description] or [No]
Following the above mentioned steps, M2CVD obtains the vulnerability assessment and description form LLMs. The formalized procedure is detailed as follows:
(2) $c_{i},n_{i}=f_{c}(p_{i}),0<i\leq M$
By analyzing the responses from the LLM, M2CVD obtains a vulnerability assessment $c$. If the LLM determines the code $p_{i}$ to be vulnerable and replies with ”Yes”, then $c_{i}=0$. Concurrently, M2CVD records the vulnerability description $n_{i}$ provided by the LLM. If the LLM determines that the code is not vulnerable, $c_{i}=1$ and $n_{i}$ is set to null.
3.2. Vulnerability Description Refinement
Through the process described above, we obtained two vulnerability assessments from the detection models and the LLM, as well as a vulnerability description from the LLM. In Phase II, we need to further refine the vulnerability assessments and descriptions from the LLM.
In this section involves a comparative analysis of the vulnerability assessments. In the case of an inconsistency between assessments, LLMs can be informed of the vulnerability assessment derived from the detection model. The second interaction enables the LLM to obtain the assessment result of the detection model that fine-tuned based on the historical data, which may enable the LLMs to regenerate its vulnerability assessment and description. The prompt for Phase II is as follows:
User: You are a senior programmer … … Code: int ff_get_wav_header(AVFormatContext *s, AVIOCon ……
LLM: [Yes][Vulnerability description] or [No]
User: Another expert has found that the code [does not] have vulnerabilities, please recheck it, and If you believe there are vulnerabilities, reply starting with ’Yes’ and briefly explain the issue; otherwise, begin with ’No’.”
LLM: [Yes][Vulnerability description] or [No]
Limiting Phase II to code fragments with divergent prediction outcomes can significantly reduce LLM inference time. M2CVD then proceeds to refine the vulnerability descriptions based on this streamlined approach:
(3) $c_{i},n_{i}=\begin{cases}c_{i},n_{i}&\text{ if }c_{i}==z_{i}\\ f_{c}(p_{i},z_{i})&\text{ else }\end{cases}$
When the assessment results from both models are consistent, the interaction with ChatGPT is terminated. This also means that the refinement process does not trigger. When the models yield inconsistent assessments, i.e., $c_{i}\neq z_{i}$, a second round of interaction is initiated using the aforementioned prompt. The values of $c_{i}$ and $n_{i}$ are updated accordingly.
3.3. Integrated Vulnerability Detection
In Phase III, M2CVD leverages the vulnerability assessments and vulnerability descriptions from the LLM to supplement the vulnerability semantics of the vulnerability code. Initially, we restructure the input dataset $L$. such that $L_{c}=(P,C,N)$. The $C$ represents ChatGPT’s assessment, while $N$ denotes the LLM’s description of the vulnerability in the code segment. After integrating the vulnerability semantics into the dataset $L_{c}=(P,C,N)$, the validation model $f_{v}$ was obtained by fine-tuning.
Algorithm 1 M2CVD
0:  $L=(P,Y)$, code model $f_{d}$,$f_{v}$,LLM:$f_{c}$
1:  Split the dataset $L$ into $p_{t},p_{v},p_{e}$
2:  Fine-tune the detection model $f_{d}$ through $p_{t}$
3:  The assessment results of detection models on the dataset were calculated: $z_{i}=f_{d}(p_{i})$
4:  for $i=1$ to $\text{len}(p_{i})$ do
5:     $c_{i},n_{i}=f_{c}(p_{i})$
6:     if $z_{i}\neq c_{i}$ then
7:        $c_{i},n_{i}=f_{c}(p_{i},z_{i})$
8:     end if
9:  end for
10:  The new data: $p^{{}^{\prime}}_{i}=p_{i}+n_{i}$
11:  Fine-tune the validation model $f_{v}$ through $p^{{}^{\prime}}_{t},p^{{}^{\prime}}_{v}$Inference phase
12:  The assessment results of validation model were calculated: $\hat{y}=f_{v}(p^{{}^{\prime}}_{e})$
13:  return  $\hat{y}$
For the fine-tuning process of the validation model, it is crucial to ensure that the training set undergoes both Phase I and II for filling in vulnerability semantics. This approach differs significantly from merely using a vulnerability label for semantic completion via the LLM. We have found that limiting the enhancement of code vulnerability semantics to only those with identified vulnerabilities in the training set can lead to overfitting in the validation model. Consequently, to adequately supplement the entire training set’s semantic content, historical data plays a pivotal role in the vulnerability semantic generation process during Phases I and II. This necessity occasionally leads ChatGPT to erroneously add semantics to codes perceived as vulnerable. Nevertheless, maintaining the same data source during both the training and inference phases, while integrating a degree of noise, has effectively increased the robustness of the validation model.
3.4. Inference Phase
In the inference phase as shown in algorithmic 1, we follow phases 1 and 2 to get the LLM’s vulnerability description $n$ and assessment $c$ for the code to be detected $p_{e}$. Finally, it was combined as the input of the validation model to complete the vulnerability detection task. The code to be detection: $p_{e}^{{}^{\prime}}=<p_{e},c,n>$ The formal procedure is articulated as follows:
(4) $\hat{y_{i}},n_{i}=f_{v}{(p_{e}^{{}^{\prime}})},0<i\leq M$
Here, $\hat{y_{i}}$ signifies the ultimate assessment result corresponding to $p_{e}$. $n_{i}$ is the vulnerability description of the code to be tested. After the M2CVD process, $n_{i}$ can be used to assist programmers to modify the vulnerability. The inference of the detection model and LLM for the code under test are performed simultaneously, and usually the inference time of LLM is greater than that of the detection model. Therefore, the inference cost is the time to call the LLM API plus the time with the validation model.
3.5. Loss Function
The loss function adopted for the code models training is the cross-entropy loss (Zhou et al., 2019), commonly used in classification problems for its effectiveness in penalizing the predicted labels and the actual labels:
(5) $H(y,\hat{y})=-y\log(\hat{y})-(1-y)\log(1-\hat{y})$
4. Experiments
Table 1. Comparison results for different models on Devign and Reveal datasets.The best result for each metric is highlighted in bold. The cells in grey represent the performance of the top-3 best methods in each metric, with darker colors representing better performance.
Dataset Devign (Zhou et al., 2019) Reveal (Chakraborty et al., 2021)
Models Accuracy Recall Precision F1 Accuracy Recall Precision F1
ChatGPT3.5-COT 49.83 32.24 33..00 30.61 63.72 26.34 30.54 27.70
Devign 56.89 52.50 64.67 57.59 87.49 31.55 36.65 33.91
SySeVR 47.85 46.06 58.81 51.66 74.33 40.07 24.94 30.74
ReGVD 61.89 48.20 60.74 53.75 90.63 14.47 64.70 23.65
CodeBERT 63.59 41.99 66.37 51.43 91.33 34.64 62.20 44.50
UniXcoder 64.82 54.48 63.53 58.94 91.90 26.31 78.94 39.47
CodeT5 65.04 54.26 64.12 58.78 88.79 38.16 43.28 40.56
M2CVD 68.11 54.50 69.51 61.09 92.12 26.75 83.56 40.56
4.1. Datesets
To evaluate the effectiveness of M2CVD, we employ two datasets from real projects :(1) Devign  (Zhou et al., 2019), and (2) Reveal (Chakraborty et al., 2021). The Devign dataset, derived from a graph-based code vulnerability detection study  (Zhou et al., 2019), stands as a dataset of function-level C/C++ source code from the well-established open-source projects QEMU and FFmpeg. Aligning with the methodology articulated by Li et al.  (Zhou et al., 2019), the partitioning of the Devign dataset adheres to a conventional 80:10:10 ratio, demarcating the bounds for training, validation, and testing data, respectively. The dataset completes the labeling of vulnerable code by a group of security researchers performing a rigorous two-stage review. In the task of software vulnerability detection, the REVEAL dataset is a representative dataset, as presented in  (Chakraborty et al., 2021). It is a further exploration of data redundancy and unrepresentative class distributions in existing datasets. As a detection code dataset, REVEAL encompasses source code extracted from two open-source forays: the Linux Debian kernel and Chromium. Similar to the real-world situation, this dataset has an imbalanced label distribution, with the number of normal code fragments much larger than the number of vulnerable ones (10:1). Similarly, in the Reveal dataset, a split ratio of 80:10:10 was set.
During the experiment, the proportion of positive and negative samples in the training set, validation set and test set is consistent with the original dataset.
4.2. Performance Metrics
In the process of evaluating the performance of the model, the proposed method includes four metrics widely recognized in the field of software testing and analysis(Zhou et al., 2019):
Precision: Denoted as the quotient of the sum of true positives and false positives and is a measure of the accuracy of instances that are identified as positive. Formally, it is defined as:
(6) $Precision=\frac{TP}{TP+FP}$
where TP and FP represent the number of true positives and false positives, respectively.
Recall: Recall evaluates the fraction of actual positives that are correctly identified and is calculated as the fraction of true positives over the sum of true positives and false negatives:
(7) $Recall=\frac{TP}{TP+FN}$
where FN signifies the number of false negatives.
F1 Score: The F1 score provides an indicator of the accuracy of the test by combining precision and recall into a single metric by taking their harmonic mean:
(8) $F1\ Score=2\times\frac{Precision\times Recall}{Precision+Recall}$
Accuracy: This metric reflects the proportion of true positives and true negatives amongst all evaluated instances, thus offering an overall measure of the model’s performance:
(9) $Accuracy=\frac{TP+TN}{TP+TN+FN+FP}$
where TN representing the number of true negatives.
4.3. Baseline Methods
In our evaluation, we compare M2CVD with seven state-of-the-art methods.
(1) ChatGPT (openAI, 2022): The GPT series models showcases the capabilities of DL in text generation and processing, albeit not specifically tailored for the domain of software vulnerability detection. ChatGPT3.5 provides the ability to abstract code vulnerabilities at a lower cost.
(2) Devign (Zhou et al., 2019): Devign is a graph-based model that uses Gated Graph Recurrent network (GGN) to represent the graph combining AST, CFG, DFG and code sequence of the input code fragment for vulnerability detection.
(3) ReGVD (Nguyen et al., 2022): ReGVD treats the problem as text classification by transforming the source code into a graph structure, using token embedding from GraphCodeBERT (Guo et al., 2020), and applying a mixture of graph-level sum and max-pooling techniques.
(4) SySeVR (Li et al., 2021b): SySeVR uses code statements, program dependencies, and program slicing as capabilities, and utilizes bidirectional recurrent neural networks to detect vulnerable code fragments.
(5) CodeBERT (Feng et al., 2020): CodeBERT use a pre-trained structure that amalgamates natural language and programming language, facilitating a broad spectrum of coding tasks, including but not limited to code understanding and generation.
(6) CodeT5 (Wang et al., 2021): CodeT5, a unified pretrained encoder-decoder Transformer model that better leverages code semantics conveyed by identifiers assigned from developers.
(7) UniXcoder (Guo et al., 2022): This is a unified code representation model that leverages a Transformer-based architecture. UniXcoder extends the capabilities of models like CodeBERT by incorporating a comprehensive understanding of code syntax and semantics, thereby enhancing the model’s performance in coding tasks such as code summarization, translation, and completion.
Experiment Environment: We implemented M2CVD in Python using Tensorflow. The experiments were performed on a machine containing a NVIDIA GeForce GTX A6000 Gpu and two Intel Xeon Gold 6226R 2.90 GHz cpus.
5. EXPERIMENTS
In this section, we conduct extensive experiments to justify our model’s superiority and analyze the reasons for its effectiveness. Specifically, we aim to answer the following research questions:
RQ1: How effective is M2CVD compared with the state-of-the art baselines on vulnerability detection?
RQ2: What are the effects of vulnerability description refinement for M2CVD?
RQ3: What are the effects of vulnerability description refinement for LLMs?
RQ4: What are the effects of different pre-trained code models and LLMs working together on vulnerability detection?
5.1. RQ1. Effectiveness of M2CVD
To answer the first question, we compare M2CVD with the seven baseline methods on the two datasets as shown in table 2. We can draw conclusions about the performance of M2CVD compared to the baselines across the evaluated datasets. Here is a comprehensive performance analysis for M2CVD:
M2CVD demonstrates a marked superiority in terms of Accuracy on both datasets. In the Devign dataset, M2CVD attains the highest Accuracy of 68.11%, the highest F1 score of 61.09% and the highest Precision outperforming all other models. This indicates that M2CVD has the most balanced performance in correctly identifying vulnerabilities without being skewed towards over-predicting (which would increase recall but decrease precision) or under-predicting (which would do the opposite).
On the Reveal dataset, M2CVD again stands out with the highest Accuracy of 92.12%. Both Precision and F1 metrics maintain the level of optimal level. These figures not only show that M2CVD maintains its high performance in different testing conditions but also that it consistently understands and predicts code vulnerabilities with high precision and recall. The performance improvement of M2CVD on the reveal dataset is much less than that of Devign, which we believe is caused by the imbalance of the Reveal dataset, and the vulnerability data is far less than the normal data. This allows ChatGPT to add far less vulnerability semantics to this dataset than to the Devign dataset. Experimental reality, ChatGPT has ultra-high recall on both datasets. After case analysis, we found that ChatGPT tends to think that the code does not have vulnerabilities when the code context is not provided.
It is worth noting that the underlying code model for M2CVD is the UniXcoder. After the multi-model collaboration process, M2CVD demonstrates superior accuracy and precision over UniXcoder, with marked improvements seen in both the Devign and Reveal datasets. Overall, the performance of M2CVD shows that M2CVD has a more balanced and higher performance in overall performance compared to UniXcoder.
Result Answer to RQ1: The performance of M2CVD shows that compared with a single model, M2CVD effectively achieves higher performance in code defect detection tasks under different experimental conditions through a collaborative mechanism.
Table 2. Model Accuracy Comparison with different M2CVD configuration
Model Accuracy
CodeBERT 63.59%
M2CVD(CodeBERT) w/o PII 65.50%
M2CVD(CodeBERT) 66.10%
UniXcoder 64.82%
M2CVD(UniXcoder) w/o PII 67.05%
M2CVD(UniXcoder) 68.11%
5.2. RQ2: Effects of vulnerability description refinement for detection performance
In this section, we elaborate on the implications of Phase II feedback within the M2CVD framework on the performance of code vulnerability detection.
For this purpose, a comparative experiment was established using the Devign dataset with its default partitioning. The configurations employed in the experiment are as follows:
a) CodeBERT: Utilizes the optimal configuration as reported in existing literature.
b) UniXcoder: Also adopts the optimal configuration as documented in existing literature.
c) M2CVD(CodeBERT) w/o PII: This configuration bypasses the M2CVD comparison process, meaning that the LLMs conduct a once assessment without incorporating feedback from the code model’s assessments. The description rendered by LLMs is amalgamated with the code, and predictions are made using the CodeBERT model.
d) M2CVD(UniXcoder) w/o PII: It omits the M2CVD comparison process. The LLM’ first vulnerability detection, when combined with the code, employs the UniXcoder model for prediction.
e) M2CVD(CodeBERT) : Use the standard M2CVD process where ChatGPT as the LLMs and CodeBERT as the code model.
f) M2CVD(UniXcoder): Also use the standard M2CVD process, with ChatGPT as the LLMs and UniXcoder as the code model.
The results from the experiment as shown in the table 3, which presents a comparison results of model accuracy under different M2CVD configurations. The experimental result evidence suggests a improvement in model accuracy when integrating Phase II feedback into the M2CVD framework. Notably, the M2CVD (CodeBERT) w/o PII outperforms the CodeBERT model by a margin of 1.91%. Similarly, the M2CVD (UniXcoder) w/o PII configuration outperforms the UniXcoder by a margin of 2.23%. The enhancements are more pronounced when the comparison includes Phase II, as observed with the M2CVD (CodeBERT) configuration, which incorporates Phase II feedback, outperforms the CodeBERT model by a margin of 2.51%. And the M2CVD (UniXcoder) configuration with Phase II integration surpasses its UniXcoder counterpart by 3.29%.
The results confirm the proposed concept. By integrating vulnerability semantics into code data, we enhance the prediction accuracy of the pre-trained code model in vulnerability detection tasks. Concurrently, these results highlight the efficacy of the vulnerability semantic refinement process within the M2CVD framework. This process significantly boosts the code model’s predictive capabilities during the final execution of the code judgment task.
Result Answer to RQ2: Experiments show that the vulnerability description refinement process of M2CVD can effectively improve the performance of code vulnerability detection.
5.3. RQ3. Effects of vulnerability description refinement for LLMs
In this section, we elaborate on the implications of Phase II feedback within the M2CVD framework on the performance of the ChatGPT model.
The aim is to determine whether informing ChatGPT of the results of the code model, after the initial assessment in M2CVD, benefits ChatGPT’s performance for code vulnerability detection. For this purpose, we set up a comparative experiment based on the Devign dataset.
The configurations employed in the experiment are as follows:
a) ChatGPT: Get ChatGPT’s vulnerability assessment interactively.
b) ChatGPT-fewshow: Get ChatGPT’s vulnerability assessment interactively. Two labeled defective code segments and one labeled non-defective code segment are selected as examples using a random selection method.
c) ALL“YES” for ChatGPT: During refinement, we informed ChatGPT of the code model evaluations, but all of them were ”YES.” This meant that we were mistakenly telling ChatGPT that every fragment of code was vulnerable.
d) ALL ”NO” for ChatGPT: During refinement, we informed ChatGPT of the evaluation of code models, but all code models judged ”NO”. This means that we are wrongly telling ChatGPT that every fragment of code is free of bugs.
e) UniXcoder for CodeLLaMa: Using the standard M2CVD process, the LLMs was chosen as CodeLLaMa.
f) UniXcoder for ChatGPT: Using the standard M2CVD process, the LLMs was chosen as ChatGPT.
where CodeLLaMa-13B is a large language models based on Llama 2. It provides excellent performance among open-source LLMs with long input contexts (Roziere et al., 2023).
Table 3. LLMs Accuracy Comparison with different configurations in Phase II
Model Accuracy
ChatGPT 45.29%
ChatGPT-fewshot 45.30%
CodeLLaMa-13B 49.57%
UniXcoder for CodeLLaMa-13B 52.59%
ALL“YES” for ChatGPT 50.04%
ALL ”NO” for ChatGPT 52.12%
UniXcoder for ChatGPT 57.61%
Based on the experimental data provided in Table 4, we can conclude that refinement step of the M2CVD framework plays a significant role in enhancing the performance of LLMs in detecting code vulnerabilities. The experiments reveal several insights: The experimental data in Table 5 underscores the impact of the initial feedback given to ChatGPT during Phase II of the M2CVD process. Firstly, ChatGPT-fewshot has almost no improvement over ChatGPT. This is due to the many types and complex forms of code defects, and a small number of instances cannot provide effective reference for large models. When ChatGPT is informed that an expert has judged a piece of code as vulnerable (”YES”), its accuracy in detecting code vulnerabilities increases from 45.29% to 50.04%. This suggests that ChatGPT benefits from additional context. Even by prompting him with insufficient information to check the data again, the detection accuracy can be improved. Similarly, if ChatGPT is consistently informed that an expert has judged the code as not vulnerable (”NO”), the model’s accuracy further improves to 52.12%. When the evaluation results of the UniXcoder model fine-tuned on the dataset are provided for ChatGPT, the accuracy of vulnerability detection rises to 57.61%. This suggests that the UniXcoder model encapsulates the dataset’s inherent logic effectively and can guide the LLM (ChatGPT) towards more accurate evaluations. On the CodeLLaMa model, we observe the same phenomenon, increasing the accuracy from 49% to 52%.
The accuracy of the basic ChatGPT model is 45.29%, while after the refinement process of M2CVD, ChatGPT shows significant performance improvement. This suggests that from a specialized code model, which carries insights from its fine-tuning process, is crucial in helping LLMs better understand and evaluate the code in question.
Result Answer to RQ3: The strategy of informing LLMs with the insights from code models that are attuned to the specific dataset logic not only improves the performance but also highlights the potential of collaborative learning systems in code vulnerability detection. This process effectively enhances the accuracy of LLMs in the code vulnerability detection task within the framework of M2CVD, and also enhances the accuracy of LLMs in adding vulnerability semantics.
5.4. RQ4. Effects of different pre-trained code models and LLMs working together
In this section, we elaborate on the impact of different code models and LLMs on the performance of M2CVD vulnerability prediction. The aim is to judge the impact of the choice of code model and LLMs during M2CVD on the final performance. To this end, we set up comparative experiments using the Devign dataset and its default partition.
The configurations employed in the experiment are as follows:
a) CodeBERT: Utilizes the optimal configuration as reported in existing literature.
b) UniXcoder: Also adopts the optimal configuration as documented in existing literature.
c) CodeLLaMa-13B+CodeBERT: Also use the standard M2CVD process, with CodeLLaMa serving as the LLMs and CodeBERT as the code model.
d) CodeLLaMa-13B+UniXcoder: Also use the standard M2CVD process, with CodeLLaMa serving as the LLMs and UniXcoder as the code model.
e) ChatGPT+CodeBERT: Also use the standard M2CVD process, with ChatGPT serving as the LLMs and CodeBERT as the code model.
f) ChatGPT+UniXcoder: Also use the standard M2CVD process, with ChatGPT serving as the LLMs and UniXcoder as the code model.
The result from the experiment presented in Table 5 provides experimental result about the impact of combining different code models and LLMs within the M2CVD framework for predicting code vulnerabilities. The standalone models, CodeBERT and UniXcoder, establish a baseline with accuracy of 63.59% and 64.82%, respectively. The combination of CodeBERT with CodeLLaMa-13B results in a slight accuracy increase, reaching 64.38%. When UniXcoder is paired with CodeLLaMa-13B, there is a more noticeable improvement, with the accuracy climbing to 64.93%. These figures serve as a benchmark to assess the added value of integrating LLMs with code models. More substantial gains in accuracy are observed when ChatGPT is introduced to the mix. ChatGPT paired with CodeBERT yields an accuracy of 66.10%, while its combination with UniXcoder tops the table at 68.11%. This indicates that the ChatGPT model significantly enhances the performance of both code models, with the ChatGPT+UniXcoder configuration proving to be the most effective partnership in this experiment.
Result Answer to RQ4: The synergy between LLMs and code models in the M2CVD framework significantly enhances the precision of detecting code vulnerabilities. Specifically, the more superior performing LLMs and code models contributes to a more pronounced accuracy improvement in the M2CVD synergy mechanism.
Table 4. Model Accuracy Comparison with different code models and LLMs working together
Model Accuracy
CodeBERT 63.59%
CodeLLaMa-13B+CodeBERT 64.38%
ChatGPT+CodeBERT 66.10%
UniXcoder 64.82%
CodeLLaMa-13B+UniXcoder 64.93%
ChatGPT+UniXcoder 68.11%
6. Case Study
In this section, we show instances of LLMs generate code vulnerability semantics. This is the core idea of the M2CVD method. Figure 2 presents a fragment of vulnerable code from the Devign dataset. This function spans over 100 lines, yet the vulnerability is concealed within just a few of them. Traditional vulnerability detection models dissect this code into tokens to learn the vulnerability features. However, these features tend to be obfuscated by a large number of no-vulnerability codes, making the learning process challenging.
Figure 2. A fragment from the Devign dataset with a code vulnerability. The vulnerability is in a few lines in this very long code.
Figure 3 provides a natural language description of the vulnerability present in this code by LLMs. This model condenses the risky elements of the code into a succinct natural language summary. This approach offers two significant advantages. Firstly, the simplification of features substantially eases the code model’s learning process, focusing on specific keywords related to the vulnerabilities. Secondly, it provides a more consistent representation of vulnerabilities. Similar vulnerabilities are expressed diversely across different code segments, leading to a vast feature space for vulnerability characteristics. Language models alleviate this issue by standardizing the same vulnerabilities into a unified natural language description, thereby streamlining the learning and identification process.
Figure 3. ChatGPT’s abstract representation of the vulnerability in this code fragment(fig.2).
7. DISCUSSION
In this section, we discuss the design of prompt in M2CVD, the reason for version selection of ChatGPT model.
Design of prompt. In the collaborative process of the M2CVD method, two sentences of intuition-based prompt are used to complete the interaction with ChatGPT. In the prompt used by M2CVD, we followed the experience provided in existing research, setting roles for LLMs and providing task contexts. Existing literature acknowledges the impact that varying prompts can have on the outcomes yielded by LLMs, with techniques such as Chain-of-Thought (Wei et al., 2022). However, the focus of this study is mainly to explore the feasibility of multi-model collaboration rather than optimization techniques of prompt, which is a concern more related to the field of prompt technology. Although the prompts employed within M2CVD may not represent the zenith of optimization, their application has resulted in a significant enhancement of performance in the code vulnerability detection tasks, underscoring the efficacy of the multi-model collaborative approach.
Version of ChatGPT. In its latest release, ChatGPT 4 offers much more powerful generation and understanding capabilities than ChatGPT 3.5. However, due to individual user usage limits, each person is currently only allowed to access the API 200 times. At the same time, it needs to pay a relatively expensive usage fee, which makes it difficult to generate the vulnerability semantics of tens of thousands of code fragments through ChatGPT 4. In comparison, ChatGPT 3.5 has a more relaxed access policy and price, but also a more powerful understanding and generation capability.
8. CONCLUSION
In this paper, we introduced M2CVD, a novel method designed to address the challenge of software vulnerability detection by harnessing the combined strengths of pre-trained code models and large language models. The M2CVD integrates the language models such as ChatGPT and code models like UniXcoder, to create a collaboration process capable of detecting vulnerabilities with high accuracy. Empirical evaluations conducted on the REVEAL and Devign datasets have demonstrated the effectiveness of M2CVD, showcasing its superior performance in detecting code vulnerabilities compared to existing benchmarks. The results of this research not only confirm the viability of M2CVD as a high-fidelity detection system but also underscore the potential of model synergy in enhancing the capabilities of automated vulnerability detection mechanisms. In essence, M2CVD demonstrates the potential to exploit the ability of different models to work together, providing a new idea for future research in automated software vulnerability detection and a scalable and effective solution for protecting software systems from changing threats.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.